Plea2. a) (7 pnts) Solve the second order homogeneous linear differential equation y - y = 0. b) (6 pnts) Without any solving, e3. a) (7 pnts) Find all eigenvalues of the matrix A = 3 -5 3 16 -6 4 11 -3 3 b) (7 pnts) Find all eigenvectors of the matrixse solve them in clear hand write . Thankyyyoouu
Answer #1

-> (2+1) { 15+1(4-1) 1% - 3(4-1) +18] =0 = (2+1) (4-x) (5+1-3) = 0 (2 + x)² (4-1) = 0 1=-2, -2, 4 Eigen values of A -2,-2.4.m 2. a) - y = 0 Let y=emu (to) be the troial solution. Auxiliary equation m² - 1 - -> 1 => m = 1 1 The solution 7 ciek & each